# Learning to Segment for Vehicle Routing Problems

### v1

@@ C.4 Parameters and Training Hyperparameters
Regarding network architecture, our encoder maps node features $\mathbf{X}\in\mathbb{R}^{n\times 25}$ for standard problems ($\mathbf{X}\in\mathbb{R}^{n\times 28}$ for VRPTWs) to node embeddings via $\mathbf{h}^{\mathrm{init}}_{i}\!=\!\text{Concat}(\mathbf{h}^{\mathrm{MLP}}_{i},\mathbf{h}^{\mathrm{POS}}_{i})\in\mathbb{R}^{2d_{h}}$, where $d_{h}=128$. They then undergo processing through $L_{\mathrm{TFM}}=2$ Transformer layers [47] with route-specific attention masks, followed by a Graph Attention Network to derive the final node embeddings $\mathbf{H}^{\mathrm{GNN}}$. The transformer implementation utilizes 2 attention heads, 0.1 dropout regularization, ReLU activation functions, layer normalization, and feedforward dimensionality of 512. Our GNN employs a transformer convolution architecture with 2 layers ($L_{\text{GNN}}=2$) and a single attention head.
@@
Figure 10: Training data construction process for the AR model. The process begins with edge modifications identified after re-optimization: deleted edges (blue and green dashed lines) and newly inserted edges (red dashed lines). Connected components are formed from these modified edges (shown in dashed circles). For each connected component, a depth-first search generates a node sequence alternating between deletion and insertion stages, with edges treated as undirected to allow operations from either endpoint. The node sequence serves as a training label for the AR model, terminated by an end token.
C.5 Training Data Construction for the AR Model
We illustrate the dataset construction process for training the AR model in Figure 10. After employing backbone solvers as lookahead solvers to re-optimize the routing problems, we identify two sets of edge modifications: deleted edges ($E_{\mathcal{R}}\setminus E_{\mathcal{R}_{+}}$), represented by blue and green dashed lines, and inserted edges ($E_{\mathcal{R}_{+}}\setminus E_{\mathcal{R}}$), represented by red dashed lines. The set $E_{\text{diff}}$ encompasses all such modified edges. Subsequently, we identify the connected components induced by $E_{\text{diff}}$, as shown within the dashed circles. For each connected component, we perform a depth-first search to generate a node sequence that alternates between connections via deleted edges (corresponding to the deletion stage) and inserted edges (corresponding to the insertion stage). During the depth-first search, all edges are treated as undirected, meaning the label for each stage (insertion or deletion) can originate from either endpoint. This sequence serves as a training label for the AR model and is terminated with a special end token to indicate the completion.
Appendix D Additional Experiments and Analysis
@@ D.1 Hyperparameter Study
D.2 Results on Realistic Routing Datasets
We further evaluate L2Seg on the CVRPLib realistic routing dataset [56, 57], adhering to the settings established in [26], which incorporates instances from CVRP Set-X [54] and the very large-scale CVRP dataset Set-XXL in the test set. The instances within CVRPLib exhibit more realistic spatial distributions (distinct from simplistic uniform or clustered patterns), greater diversity, and better representation of real-world logistical challenges. For this evaluation, we employ models trained on synthetic small-capacity CVRP2k and CVRP5k datasets and zero-shot transfer them to CVRPLib. Time constraints of 240s and 600s are implemented for L2Seg during testing. Additional methodological details are provided in Appendix C. As demonstrated in Table 10, LNS augmented with L2Seg-SYN surpasses all other learning-based methods in performance. Significantly, the computational time required by LNS+L2Seg-SYN (600s) is substantially less than that of the previously best-performing learning-based model, UDC-$\boldsymbol{x}_{250}$. These results further substantiate L2Seg’s exceptional generalizability across varied problem distributions.
@@ D.2 Results on Realistic Routing Datasets
Dataset, $N\in$ UDC-$\boldsymbol{x}_{2}$ UDC-$\boldsymbol{x}_{250}$ LNS+L2Seg-SYN (240s) LNS+L2Seg-SYN (600s)
@@
D.3 Standard Deviation Comparison
@@
D.4 Case Study: Comparison of Predictions of Three L2Seg Approaches
@@
We present a case study on a small-capacity CVRP1k instance to analyze model prediction behavior. Since the learned model ultimately predicts on two adjacent routes, we visualize unstable edge predictions (red dashed lines) for two such routes using L2Seg-SYN, L2Seg-NAR, and L2Seg-AR in Figure 12. L2Seg-SYN demonstrates selective prediction behavior, avoiding boundary edges while targeting specific unstable edges within route interiors—a pattern consistent with our observations in Appendix A.1.1. L2Seg-NAR successfully identifies unstable regions (route interiors) but lacks discrimination, predicting nearly all edges within these regions as unstable without capturing local dependencies. L2Seg-AR exhibits selective prediction within regions but fails to properly identify unstable regions, as many predictions occur at boundaries. These results provide insight into L2Seg-SYN’s hybrid approach: the NAR component first identifies unstable regions, while the AR component leverages local information to make accurate predictions within each identified region.
Appendix E Code Availability and Licenses
Our complete implementation will be made publicly available on GitHub under the MIT License upon acceptance of this paper. All assets utilized in our research are properly attributed, and we have included licensing information as below:
  1. 1.
    MIT License: HGS [8], UDC [26], SIL [5], L2R [42], L2D [4]
  2. 2.
    Available for academic use: LKH-3 [7],
    CVRPLib (http://vrp.galgos.inf.puc-rio.br/index.php/en/)

### v2

@@ D.3 Parameters and Training Hyperparameters
Regarding network architecture, our encoder maps node features $\mathbf{X}\in\mathbb{R}^{n\times 25}$ for standard problems ($\mathbf{X}\in\mathbb{R}^{n\times 28}$ for VRPTWs) to node embeddings via $\mathbf{h}^{\mathrm{init}}_{i}\!=\!\text{Concat}(\mathbf{h}^{\mathrm{MLP}}_{i},\mathbf{h}^{\mathrm{POS}}_{i})\in\mathbb{R}^{2d_{h}}$, where $d_{h}=128$. They then undergo processing through $L_{\mathrm{TFM}}=2$ Transformer layers (Vaswani, 2017) with route-specific attention masks, followed by a Graph Attention Network to derive the final node embeddings $\mathbf{H}^{\mathrm{GNN}}$. The transformer implementation utilizes 2 attention heads, 0.1 dropout regularization, ReLU activation functions, layer normalization, and feedforward dimensionality of 512. Our GNN employs a transformer convolution architecture with 2 layers ($L_{\text{GNN}}=2$) and a single attention head.
@@
D.4 Instance Generation
In general, we generate all training and test instances following established methodologies: Zheng et al. (Zheng et al., 2024) for CVRP and Solomon (Solomon, 1987) for VRPTW. Specifically, For small-capacity CVRPs, nodes are uniformly distributed within the $[0,1]$ square, with integer demands ranging from 1 to 9 (inclusive). Vehicle capacities are set to $C=200$, 300300, and $300$ for problem sizes 1k, 2k, and 5k, respectively. For large-capacity CVRPs, we maintain identical configurations except for increased vehicle capacities of $C=500$ and $1000$ for CVRP1k and CVRP5k, respectively. For VRPTWs, we adopt the same spatial distribution, demand structure, and capacity constraints as the small-capacity CVRPs. Service times are uniformly set to 0.2 time units for each customer and 0 for the depot. Time windows are generated according to the methodology outlined in Solomon (Solomon, 1987).
  • Validation: 30 instances per problem configuration
  • Testing: For small-capacity CVRPs, we utilize the 1,000 test instances from Zheng et al. (Zheng et al., 2024); for large-capacity CVRPs and VRPTWs, we evaluate on 100 instances sampled from the same distribution as the training data
Appendix E Additional Experiments and Analysis
@@ E.1 Hyperparameter Study
E.2 Results on Realistic Routing Datasets
We further evaluate L2Seg on the CVRPLib realistic routing dataset (Uchoa et al., 2017, Arnold et al., 2019), adhering to the settings established in Zheng et al. (2024), which incorporates instances from CVRP Set-X [54] and the very large-scale CVRP dataset Set-XXL in the test set. The instances within CVRPLib exhibit more realistic spatial distributions (distinct from simplistic uniform or clustered patterns), greater diversity, and better representation of real-world logistical challenges. For this evaluation, we employ models trained on synthetic small-capacity CVRP2k and CVRP5k datasets and zero-shot transfer them to CVRPLib. Time constraints of 240s and 600s are implemented for L2Seg during testing. Additional methodological details are provided in Appendix D. As demonstrated in Table 10, LNS augmented with L2Seg-SYN surpasses all other learning-based methods in performance. Significantly, the computational time required by LNS+L2Seg-SYN (600s) is substantially less than that of the previously best-performing learning-based model, UDC-$\boldsymbol{x}_{250}$. These results further substantiate L2Seg’s exceptional generalizability across varied problem distributions.
@@ E.2 Results on Realistic Routing Datasets
Dataset, $N\in$ $\boldsymbol{x}_{2} UDC$-$\boldsymbol{x}_{250}$ LNS+L2Seg-SYN (240s) LNS+L2Seg-SYN (600s)
@@
E.3 Results on Clustered CVRP and Heterogeneous-demand CVRP
Table 11: Results on clustered CVRP and heterogeneous-demand CVRP. We present gains to the backbone solver LNS and the performance of LKH-3 for reference.
Methods Clustered CVRP2k Clustered CVRP5k
[…]
LKH-3 (Helsgaun, 2017) (for reference) 42.06 - 150s 62.33 - 240s
LNS (Shaw, 1998) 41.54 0.00% 150s 61.42 0.00% 240s
L2Seg-SYN-LNS (zero-shot transfer) 41.03 1.23% 150s 60.87 0.90% 240s
L2Seg-SYN-LNS 40.73 1.95% 150s 60.11 2.13% 240s
Methods Hetero-demand CVRP2k Hetero-demand CVRP5k
[…]
LKH-3 (Helsgaun, 2017) (for reference) 46.02 - 150s 65.89 - 240s
LNS (Shaw, 1998) 45.77 0.00% 150s 64.81 0.00% 240s
L2Seg-SYN-LNS (zero-shot transfer) 44.35 3.10% 150s 64.28 0.82% 240s
L2Seg-SYN-LNS 44.15 3.54% 150s 64.15 1.02% 240s
To demonstrate L2Seg’s robustness across diverse and more realistic scenarios beyond uniform distributions, we provide in-distribution and zero‑shot generalization evaluation of our L2Seg on instances with different customer and demand distributions.
Following Li et al. (2021), we generate clustered CVRP instances with 7 clusters. For heterogeneous-demand scenarios, we employ a skewed distribution where high and low demands ($d\in\{1,2,8,9\}$) occur with probability 0.2 each, while others ($d\in\{3,4,5,6,7\}$) occur with probability 0.04 each. All experiments use LNS as the backbone solver, with LKH-3 included for reference.
Table 11 presents the comprehensive results. L2Seg demonstrates consistent improvements across all settings: zero-shot transfer achieves 1.23% to 3.10% gains over LNS, while in-distribution testing reaches 1.02% to 3.54% improvements depending on problem size and variant. These experiments demonstrate that L2Seg maintains consistent improvements across diverse real-world conditions, from uniform spatial layouts to clustered distributions and heterogeneous demands.
E.4 Standard Deviation Comparison
@@
E.5 Case Study: Comparison of Predictions of Three L2Seg Approaches
@@
We present a case study on a small-capacity CVRP1k instance to analyze model prediction behavior. Since the learned model ultimately predicts on two adjacent routes, we visualize unstable edge predictions (red dashed lines) for two such routes using L2Seg-SYN, L2Seg-NAR, and L2Seg-AR in Figure 12. L2Seg-SYN demonstrates selective prediction behavior, avoiding boundary edges while targeting specific unstable edges within route interiors—a pattern consistent with our observations in Appendix B.1.1. L2Seg-NAR successfully identifies unstable regions (route interiors) but lacks discrimination, predicting nearly all edges within these regions as unstable without capturing local dependencies. L2Seg-AR exhibits selective prediction within regions but fails to properly identify unstable regions, as many predictions occur at boundaries. These results provide insight into L2Seg-SYN’s hybrid approach: the NAR component first identifies unstable regions, while the AR component leverages local information to make accurate predictions within each identified region.
E.6 Unstable and Stable Edges Convergence
We conducted experiments measuring overlapping predicted edges between adjacent iterations over the first 10 rounds, revealing interesting dynamics: The overlap of predicted unstable edges increases from 28% to 54%, while stable edge overlap increases from 47% to 69% across iterations, shown in the Table 13. This indicates gradual but not rapid convergence, allowing our method to continuously explore new regions for re-optimization rather than getting trapped in fixed segments.
Table 13: Unstable and stable edges convergence at the first 10 iterations
Round # 1 2 3 5 7 9
Unstable Edge Overlapping Percentage 28.2% 33.5% 41.2% 49.2% 48.8% 54.1%
Stable Edge Overlapping Percentage 47.2% 58.2% 60.5% 64.7% 67.3% 69.4%
Avg Segment Length 2.45 2.57 2.44 3.04 2.87 2.73
@@ Appendix F Broader Impacts
Appendix G Large Language Models Usage
We used LLMs to assist with manuscript revision. After completing the initial draft without LLM assistance, we consulted LLMs for suggestions on improving specific text passages. All LLM-generated advice was carefully reviewed to ensure accuracy before incorporation. LLMs were not used for research tasks or any purpose beyond text refinement.